### v1

@@ Abstract
Multi-exit network is a promising architecture for efficient model inference by sharing backbone networks and weights among multiple exits. However, the gradient conflict of the shared weights results in sub-optimal accuracy. This paper introduces Deep Feature Surgery (DFS), which consists of feature partitioning and feature referencing approaches to resolve gradient conflict issues during the training of multi-exit networks. The feature partitioning separates shared features along the depth axis among all exits to alleviate gradient conflict while simultaneously promoting joint optimization for each exit. Subsequently, feature referencing enhances multi-scale features for distinct exits across varying depths to improve the model accuracy. Furthermore, DFS reduces the training operations with the reduced complexity of backpropagation. Experimental results on Cifar100 and ImageNet datasets exhibit that DFS provides up to a 50.00% reduction in training time and attains up to a 6.94% enhancement in accuracy when contrasted with baseline methods across diverse models and tasks. Budgeted batch classification evaluation on MSDNet demonstrates that DFS uses about $\mathbf{2}\boldsymbol{\times}$ fewer average FLOPs per image to achieve the same classification accuracy as baseline methods on Cifar100.
@@ 1 Introduction
  • We propose DFS, which consists of feature partitioning and feature referencing approaches to address the gradient conflict issue of training multi-exit networks.
@@ 2 Related work
Training multi-exit networks is challenging since the gradients from multiple exits hold conflicts, which usually lead to sub-optimal model accuracy [34]. Many efforts have been made to solve the gradient conflict issue as follows.
@@ 3 Preliminary: Gradient Conflict
$c_{i}=\pi_{i}(f_{i},w_{c_{i}}),~{}~{}s.t.\begin{cases}f_{i}=\varphi_{i}(X,W_{i}),\\ i=1,2,\cdots,L\end{cases}$ (1)
@@
$W_{i}\cap W_{j}=\{w_{1},w_{2},\cdots,w_{\min(i,j)}\},~{}~{}~{}s.t.\begin{cases}i=1,2,\cdots,L\\ j=1,2,\cdots,L\end{cases}$ (2)
@@
$\begin{split}g_{w_{i}}=\sum_{k=i}^{L}g^{k}_{w_{i}},~{}~{}s.t.~{}~{}g^{k}_{w_{i}}=\frac{\partial CE(c_{k},Y)}{\partial w_{i}}\end{split}$ (4)
@@ 4 Deep Feature Surgery
DFS addresses the gradient conflict by creating both individual and shared parameter updates. DFS conducts feature partitioning and feature referencing on the feature tensors to ensure the desirable model accuracy of different exits and reduce the backward operations in multi-exit model training.
@@ 4.1 Feature Partitioning
$w_{i}=[w_{i}^{+},w_{i}^{-}]_{\beta},~{}s.t.\begin{cases}w_{i}^{+}=w_{i}[:,:\lfloor\beta C_{out}\rceil,:,:]\\ w_{i}^{-}=w_{i}[:,\lfloor\beta C_{out}\rceil:,:,:]\end{cases}$ (5)
@@
$\begin{split}c_{i}&=\pi_{i}(f_{i}^{-},w_{c_{i}}),~{}~{}~{}s.t.\begin{cases}f_{i}^{-}=\varphi_{i}(X,\{w_{1}^{+},w_{2}^{+},\cdots,w_{i-1}^{+},w_{i}^{-}\})\\ i=1,2,\cdots,L-1.\end{cases}\end{split}$ (6)
@@
$g_{w_{i}}=[\sum_{k=i+1}^{L}g_{w_{i}^{+}}^{k},g^{i}_{w_{i}^{-}}]_{\beta},~{}~{}s.t.~{}g_{w_{i}^{+}}^{k}=\frac{\partial CE(c_{k},Y)}{\partial w_{i}^{+}},~{}g^{i}_{w_{i}^{-}}=\frac{\partial CE(c_{i},Y)}{\partial w_{i}^{-}}.$ (7)
@@
Weight partitioning with a variable is hard to achieve due to the complexity of the weight tensor management. Instead of partitioning the weight tensor, we partition the features for easy deployment in practice. As shown in Fig. 2, taking the first layer as an example, the output feature $f_{1}=\varphi_{1}(X,w_{1})$ of the first layer is split into $f_{1}^{+}=\varphi_{1}(X,w_{1}^{+})$ and $f_{1}^{-}=\varphi_{1}(X,w_{1}^{-})$, which is corresponding to weight partitioning in Eq. 5. The corresponding feature partitions to Eq. 6 is as follows:
$\begin{cases}f_{i}^{-}=\varphi_{i}(X,\{w_{1}^{+},w_{2}^{+},\cdots,w_{i-1}^{+},w_{i}^{-}\})\\ f_{i}^{+}=\varphi_{i}(X,\{w_{1}^{+},w_{2}^{+},\cdots,w_{i-1}^{+},w_{i}^{+}\})\end{cases},~{}~{}s.t.~{}i=1,2,\cdots,L-1.$ (8)
@@ 4.2 Feature Referencing
Feature partitioning alleviates the gradient conflict in multi-exit network training but reduces the number of features for weight training. For example, $f_{i}^{+}$ is not available for task $c_{i}$ in Eq. 6. The reduced number of features, e.g., $f_{i}^{+}$, can easily result in the under-fitting of tasks and further affect model accuracy. We introduce feature referencing to address this issue. Feature referencing reuses these features across tasks to benefit the task performance. Based on feature referencing, task $c_{i}$ and gradients $g_{w_{i}}$ of all tasks with respect to weight $w_{i}$ are defined as follows.
$\begin{cases}c_{i}&=\pi_{i}([f_{i}^{+},f_{i}^{-}]_{\beta},w_{c_{i}})\\ f_{i}^{+}&=\varphi_{i}([f_{i-1}^{+},f_{i-1}^{-}]_{\beta},w_{i}^{+})\\ f_{i}^{-}&=\varphi_{i}([f_{i-1}^{+},f_{i-1}^{-}]_{\beta},w_{i}^{-})\end{cases},~{}~{}s.t.~{}\begin{cases}f_{0}=X,\\ i=1,2,\cdots,L-1\end{cases}$ (9)
@@ 5 Training Efficiency Improvement
$\small\begin{split}f_{i}&=[f_{i}^{+},f_{i}^{-}]_{\beta}=[\varphi_{i}(f_{i-1}\cdot w_{1}^{+}),\varphi_{i}(f_{i-1}\cdot w_{1}^{-})]_{\beta}\\ &~{}s.t.~{}f_{0}=X,i=1,2,\cdots,L.\end{split}$ (12)
@@
$\small\begin{split}1-\frac{10LN^{3}}{(12L-2)N^{3}}\in[0\%,16.67\%),~{}s.t.~{}L\in[1,+\infty).\end{split}$ (14)
@@ 6.1 Model Accuracy Evaluation
Figure 3: Accuracy of budgeted batch classification. The X-axis is the average computational budget per image for MSDNet model on the Cifar100. The Y-axis is the achieved top-1 model accuracy.
@@ 6.2 Training Efficiency Validation
We adopt the VGG7-64 in [6] with 6 convolution layers and train it on Cifar100 [15] dataset, and VGG16 [30] with 13 convolution layers and train it on ImageNet [16] dataset to evaluate the training efficiency improvement from DFS. Both models employ a Batch Normalization (BN) [14] layer after each convolution layer and the early classifiers are added after the BNs. Each classifier only contains a Global Average Pooling (GAP) [20] layer and a Fully-Connected (FC) layer for classification. The baseline methods include DSN [17] and BYOT [35]. DSN applies the vanilla training Sec. 5 which optimizes all losses with the same weight and updates the trainable parameters. BYOT introduces self-distillation supervision based on DSN, which employs the outputs and logits of the final exit to teach the early tasks. We use the latency for training one batch of samples and Frames Per Second (FPS) on the same devices to evaluate the training efficiency. The overall training latency is divided into forward latency, backward latency, and other latency for detailed comparison. The other latency includes data loading, loss computation, parameter update, etc. The lower latency and higher FPS stand for a higher training efficiency. For a fair comparison, the results are collected with different batch sizes. We train VGG7-64 on CPU (2$\times$ 20-core Intel Xeon Gold 6248) and VGG16 on GPU (1$\times$ NVIDIA RTX A6000), both are trained with 100 batches of data inputs.

### v2

@@ Abstract
Multi-exit network is a promising architecture for efficient model inference by sharing backbone networks and weights among multiple exits. However, the gradient conflict of the shared weights results in sub-optimal accuracy. This paper introduces Deep Feature Surgery (DFS), which consists of feature partitioning and feature referencing approaches to resolve gradient conflict issues during the training of multi-exit networks. The feature partitioning separates shared features along the depth axis among all exits to alleviate gradient conflict while simultaneously promoting joint optimization for each exit. Subsequently, feature referencing enhances multi-scale features for distinct exits across varying depths to improve the model accuracy. Furthermore, DFS reduces the training operations with the reduced complexity of backpropagation. Experimental results on Cifar100 and ImageNet datasets exhibit that DFS provides up to a 50.00% reduction in training time and attains up to a 6.94% enhancement in accuracy when contrasted with baseline methods across diverse models and tasks. Budgeted batch classification evaluation on MSDNet demonstrates that DFS uses about $\mathbf{2}\boldsymbol{\times}$ fewer average FLOPs per image to achieve the same classification accuracy as baseline methods on Cifar100. The code is available at https://github.com/GongCheng1919/dfs.
@@ 1 Introduction
  • We propose DFS, which consists of feature partitioning and feature referencing to address the gradient conflict issue of training multi-exit networks.
@@ 2 Related work
Training multi-exit networks is challenging since the gradients from multiple exits hold conflicts, which usually lead to sub-optimal model accuracy [34].
@@ 3 Preliminary: Gradient Conflict
$\small c_{i}=\pi_{i}(f_{i},w_{c_{i}}),~{}~{}s.t.\begin{cases}f_{i}=\varphi_{i}(X,W_{i}),\\ i=1,2,\cdots,L\end{cases}$ (1)
@@
$\small W_{i}\cap W_{j}=\{w_{1},w_{2},\cdots,w_{\min(i,j)}\},~{}~{}~{}s.t.\begin{cases}i=1,2,\cdots,L\\ j=1,2,\cdots,L\end{cases}$ (2)
@@
$\small g_{w_{i}}=\sum_{k=i}^{L}g^{k}_{w_{i}},~{}~{}s.t.~{}~{}g^{k}_{w_{i}}=\frac{\partial CE(c_{k},Y)}{\partial w_{i}}$ (4)
@@ 4 Deep Feature Surgery
DFS addresses the gradient conflict by creating both individual and shared parameter updates. DFS conducts feature partitioning and feature referencing on the feature tensors to ensure the desirable model accuracy of different exits.
@@ 4.1 Feature Partitioning
$\small w_{i}=[w_{i}^{+},w_{i}^{-}]_{\beta},~{}s.t.\begin{cases}w_{i}^{+}=w_{i}[:,:\lfloor\beta C_{out}\rceil,:,:]\\ w_{i}^{-}=w_{i}[:,\lfloor\beta C_{out}\rceil:,:,:]\end{cases}$ (5)
@@
$\small c_{i}=\pi_{i}(f_{i}^{-},w_{c_{i}}),~{}~{}~{}s.t.\begin{cases}f_{i}^{-}=\varphi_{i}(X,\{w_{1}^{+},w_{2}^{+},\cdots,w_{i-1}^{+},w_{i}^{-}\})\\ i=1,2,\cdots,L-1.\end{cases}$ (6)
@@
$\small g_{w_{i}}=[\sum_{k=i+1}^{L}g_{w_{i}^{+}}^{k},g^{i}_{w_{i}^{-}}]_{\beta},~{}~{}s.t.~{}g_{w_{i}^{+}}^{k}=\frac{\partial CE(c_{k},Y)}{\partial w_{i}^{+}},~{}g^{i}_{w_{i}^{-}}=\frac{\partial CE(c_{i},Y)}{\partial w_{i}^{-}}.$ (7)
@@
Instead of partitioning the weight tensor, we partition the features for easy deployment in practice. As shown in Fig. 2, taking the first layer as an example, the output feature $f_{1}=\varphi_{1}(X,w_{1})$ of the first layer is split into $f_{1}^{+}=\varphi_{1}(X,w_{1}^{+})$ and $f_{1}^{-}=\varphi_{1}(X,w_{1}^{-})$, which is corresponding to weight partitioning in Eq. 5. The corresponding feature partitions to Eq. 6 is as follows:
$\small\begin{cases}f_{i}^{-}=\varphi_{i}(X,\{w_{1}^{+},w_{2}^{+},\cdots,w_{i-1}^{+},w_{i}^{-}\})\\ f_{i}^{+}=\varphi_{i}(X,\{w_{1}^{+},w_{2}^{+},\cdots,w_{i-1}^{+},w_{i}^{+}\})\end{cases},~{}~{}s.t.~{}i=1,2,\cdots,L-1.$ (8)
@@ 4.2 Feature Referencing
Feature partitioning alleviates the gradient conflict in multi-exit network training but reduces the number of features for weight training. For example, $f_{i}^{+}$ is not available for task $c_{i}$ in Eq. 6. The reduced number of features, e.g., $f_{i}^{+}$, can easily result in the under-fitting of tasks. We introduce feature referencing to address this issue. Feature referencing reuses these features across tasks to benefit the task performance. Based on feature referencing, task $c_{i}$ and gradients $g_{w_{i}}$ of all tasks with respect to weight $w_{i}$ are defined as follows.
$\small\begin{cases}c_{i}&=\pi_{i}([f_{i}^{+},f_{i}^{-}]_{\beta},w_{c_{i}})\\ f_{i}^{+}&=\varphi_{i}([f_{i-1}^{+},f_{i-1}^{-}]_{\beta},w_{i}^{+})\\ f_{i}^{-}&=\varphi_{i}([f_{i-1}^{+},f_{i-1}^{-}]_{\beta},w_{i}^{-})\end{cases},~{}~{}s.t.~{}\begin{cases}f_{0}=X,\\ i=1,2,\cdots,L-1\end{cases}$ (9)
@@ 5 Training Efficiency Improvement
$\small f_{i}=[f_{i}^{+},f_{i}^{-}]_{\beta}=[\varphi_{i}(f_{i-1}\cdot w_{1}^{+}),\varphi_{i}(f_{i-1}\cdot w_{1}^{-})]_{\beta}~{}s.t.~{}f_{0}=X,i=1,2,\cdots,L.$ (12)
@@
$\small 1-\frac{10LN^{3}}{(12L-2)N^{3}}\in[0\%,16.67\%),~{}s.t.~{}L\in[1,+\infty).$ (14)
@@ 6.1 Model Accuracy Evaluation
Figure 3: Accuracy of budgeted batch classification. The X-axis is the average computational budget per image for MSDNet model on the Cifar100, and Y-axis top-1 accuracy.
@@ 6.2 Training Efficiency Validation
We adopt the VGG7-64 in [6] with 6 convolution layers and train it on Cifar100 [15] dataset, and VGG16 [30] with 13 convolution layers and train it on ImageNet [16] dataset to evaluate the training efficiency improvement from DFS. Both models employ a Batch Normalization (BN) [14] layer after each convolution layer and the early classifiers are added after the BNs. Each classifier only contains a Global Average Pooling (GAP) [20] layer and a Fully-Connected (FC) layer for classification. The baseline methods include DSN [17] and BYOT [35]. DSN applies the vanilla training in Sec. 5 which optimizes all losses with the same weight and updates the trainable parameters. BYOT introduces self-distillation supervision based on DSN, which employs the outputs and logits of the final exit to teach the early tasks. We use the latency for training one batch of samples and Frames Per Second (FPS) on the same devices to evaluate the training efficiency. The overall training latency is divided into forward latency, backward latency, and other latency for detailed comparison. The other latency includes data loading, loss computation, parameter update, etc. The lower latency and higher FPS stand for a higher training efficiency. For a fair comparison, the results are collected with different batch sizes. We train VGG7-64 on CPU (2$\times$ 20-core Intel Xeon Gold 6248) and VGG16 on GPU (1$\times$ NVIDIA RTX A6000), both are trained with 100 batches of data inputs.